# Censored extreme value estimation

### v1

@@ 4.2 Applications
which appears in Theorem 4.1 and Theorem 4.2.
Thus, let $F\in\mathcal{D}(G_{\gamma})$ and assume for convenience that $\tau_{F}>0.$ By the remark below Theorem 4.2, $g(x)=x^{\gamma_{+}},$ where $\gamma_{+}=\max(\gamma,0)$. Moreover, by Potter’s bounds, for arbitrary $\delta_{1},\delta_{2}>0$, there exists $T=T(\delta_{1},\delta_{2})$ such that for $t\geq T$
$g_{1}(x):=\max\big{(}(1-\delta_{1})x^{\gamma_{+}-\delta_{2}}),1\big{)}\leq f(x,t)\leq(1+\delta_{1})x^{\gamma_{+}+\delta_{2}}=:g_{2}(x),$
@@
Next, if $\gamma>0$ then we can use the second order Condition 3. To derive an analogous condition in case $\gamma\leq 0,$ recall that $F\in\mathcal{D}(G_{\gamma})$ is equivalent to
$\lim_{t\to\infty}\frac{U_{F}(tx)-U_{F}(t)}{a(t)}=h_{\gamma}(x),\quad x>0,$ (4.12)
@@
$\lim_{t\to\infty}\frac{U_{F}(tx)/U_{F}(t)-x^{\gamma_{+}}}{a(t)/U_{F}(t)}=h_{\gamma}(x),$ (4.13)
@@
$\tilde{g}(x)=x^{\gamma_{+}}h_{\hat{\rho}}(x)$ (4.14)
with $\hat{\rho}=\rho I(\gamma>0)+\gamma I(\gamma\leq 0)$ and $\hat{a}$ equal to $a$ in Condition 3 for $\gamma>0$ and $a/U$ in (4.13) for $\gamma\leq 0,$ respectively. Next, the conditions of Lemma 4.6, [34], are satisfied under Condition 3 for $\gamma>0$ and (4.13) for $\gamma\leq 0,$ respectively, and thus in (4.9) we can set $\hat{g}(x)=Kx^{\gamma_{+}+\delta}$ for some $K>0,$ $\delta>0$ and $t\geq T.$ Finally, since $\hat{a}$ is regularly varying in our case, it suffices to require $\sqrt{k}\hat{a}(n/k)\to\lambda$ instead of $\sqrt{k}\hat{a}(\xi_{n-k,n})\stackrel{{\scriptstyle P}}{{\to}}\lambda$ by Smirnov’s lemma (see, e.g., Lemma 2.2.3 in [15] or (A.43)).
@@
    Let $F\in\mathcal{D}(G_{\gamma})$ with $\gamma\in\amsmathbb{R}$ and $\tau_{F}>0$, and $\vartheta:[1,\infty)\to\amsmathbb{R}$ satisfy $\amsmathbb{E}\big{[}(\bar{\vartheta}(\xi^{\gamma_{+}+\delta}))^{1+\delta}\big{]}<\infty$ for some $\delta>0,$
    and $\xi\sim\text{Pareto}(1).$ Assume $\mu(t),$ $k$ and $\xi_{1},\xi_{2},\ldots$ are as in Theorem 4.1. Then $R_{k,n}$ is consistent for $\amsmathbb{E}\big{[}\vartheta(\xi^{\gamma_{+}})\big{]}.$ If, moreover, $\amsmathbb{E}\big{[}(\bar{\vartheta}(\xi^{\gamma_{+}+\delta}))^{2+\delta}\big{]}<\infty$ for some $\delta>0,$ then
@@
    Assume additionally Condition 3 for $\gamma>0,$ $\sqrt{k}\hat{a}(n/k)\to\lambda$ with finite $\lambda$ for $\hat{a}$ defined above and almost everywhere continuity of $\vartheta^{\prime}$ if $\lambda\neq 0.$ Finally, assume $\amsmathbb{E}\big{[}\xi^{\gamma_{+}+\delta}\bar{\vartheta^{\prime}}(\xi^{\gamma_{+}+\delta})\big{]}<\infty$.
    Then
@@
    where $C=\amsmathbb{E}\big{[}\tilde{g}(\xi)\vartheta^{\prime}(\xi^{\gamma_{+}})\big{]}$ and $\tilde{g}$ is defined by (4.14).
Remark 4.7.
In the above, it is always possible to find an intermediate sequence $k=k_{n}$ such that $\sqrt{k}\hat{a}(n/k)\to\lambda$ is satisfied. Indeed, if $\gamma>0$ then $\hat{a}(t)\to 0$ as $t\to\infty$ by definition (see Condition 3), and if $\gamma\leq 0$ then $\hat{a}(t)\to 0$ as $t\to\infty$ by Lemma 1.2.9, [15].
Note again that to estimate the extreme value index, Corollary 4.6 is not enough since for negative $\gamma$, $R_{k,n}$ is a consistent estimator of $\vartheta(1)$, which itself does not depend on $\gamma.$ Therefore, we now consider another function $f(x,t).$
As a starting point, note that it follows from (4.12) that
$\lim_{t\to\infty}\frac{\log U_{F}(tx)-\log U_{F}(t)}{a(t)/U_{F}(t)}=h_{\gamma_{-}}(x)$ (4.15)
with $\gamma_{-}=\min(\gamma,0)$ and the same $a(t)$ as in (4.12), see, e.g., formula (3.5.4) in [15]. The limit (4.15) implies that $\log U_{F}(t)$ is extended regularly varying with index $\gamma_{-},$ see the properties of such functions in section B.2, [15]. The relation (4.15) has previously been used to establish the asymptotic behavior of the moment estimator of EVI, see [17]. Thus, let us consider the candidate function
@@
$x\mapsto\gamma^{-1}\big{(}\max[(1\mp\delta)x^{\gamma\mp\delta}-(1\pm\delta),0]\big{)}$ (4.16)
@@
Next, the second order condition for $\log U_{F}(t)$ reads as follows
$\lim_{t\to\infty}\frac{\frac{\log U_{F}(tx)-\log U_{F}(t)}{a(t)/U_{F}(t)}-h_{\gamma_{-}}(x)}{Q(t)}=\int_{1}^{x}s^{\gamma_{-}-1}\int_{1}^{s}u^{\rho^{\prime}-1}\mathop{}\!\mathrm{d}u\mathop{}\!\mathrm{d}s=:H_{\gamma_{-},\rho^{\prime}}(x),$ (4.17)
which for the cases $\gamma=0$ and $\rho=0$ is understood to be equal to the limit of (4.18) as $\gamma\to 0$ or $\rho\to 0,$ respectively. Note that (4.17) follows from the standard extreme value theory second order condition
$\lim_{t\to\infty}\frac{\frac{U_{F}(tx)-U_{F}(t)}{a(t)}-h_{\gamma}(x)}{A(t)}=H_{\gamma,\rho}(x),$ (4.19)
@@
$\left|\frac{\log U_{F}(tx)-\log U_{F}(t)}{a(t)/U_{F}(t)}-h_{\gamma_{-}}(x)\right|\leq K|Q(t)|\,x^{\gamma_{-}+\delta},\quad x\geq 1,t\geq T.$ (4.20)
Clearly, $\delta$ in (4.20) can be chosen such that $\gamma+\delta<0$ for negative $\gamma.$ Finally, $Q(t)$ appearing in (4.17) is regularly varying with index $\rho^{\prime}$ by Theorem B.3.1, [15], therefore as above we can require $\sqrt{k}Q(n/k)\to\lambda$ instead of $\sqrt{k}Q(\xi_{n-k,n})\to\lambda.$
@@
for some function $\vartheta$ and denote
@@
    Let $F\in\mathcal{D}(G_{\gamma})$ with $\gamma\in\amsmathbb{R}$ and $\tau_{F}>0.$ Let $\vartheta:[0,\infty)\to\amsmathbb{R},$ and if $\gamma\geq 0,$ satisfy $|\vartheta(x)|\leq A\max(x^{c},1),$ $x\geq 0,$ for some $A>0$ and $c>0.$ Assume $k$ and $\xi_{1},\xi_{2},\ldots$ are as in Theorem 4.1. Then $T_{k,n}$ is consistent for $\amsmathbb{E}\big{[}\vartheta(h_{\gamma_{-}}(\xi))\big{]}$ and
@@
    Assume additionally (4.17), $\sqrt{k}Q(n/k)\to\lambda$ with finite $\lambda$ for $Q$ defined in (4.17), $|\vartheta^{\prime}(0)|<\infty,$ and almost everywhere continuity of $\vartheta^{\prime}$ if $\lambda\neq 0.$ Finally, if $\gamma\geq 0,$ assume $|\vartheta^{\prime}(x)|\leq A\max(x^{c},1),$ $x\geq 0,$ for some $A>0$ and $c>0.$ Then
@@
4.3 Extension of the results of Section 3
In this section, we extend the results of Section 3 using ideas from Section 4.1. Let us use the same notation as in Section 3. Crucially, unlike in Section 3, we do not assume that $F,$ $G$ and $H$ are regularly varying distribution functions.
Let us consider a sequence $\xi,\xi_{1},\xi_{2},\ldots$ of independent Pareto$(1)$ random variables such that a.s. $Z_{i}=U_{H}(\xi_{i}),$ $i\geq 1,$ and a function $f(x,t)$ satisfying Definition 4. Let $\vartheta:[0,\infty)\to\amsmathbb{R}$ be a measurable function as in Section 4.1. The following extreme Kaplan-Meier integral is of our main interest:
where $\amsmathbb{F}^{\xi}_{k,n}$ is the extreme Kaplan-Meier estimator built by the sequence $\{\xi_{i}\},$ i.e.

### v5

@@ Appendix A Applications of generalized residual estimators
which appears in Theorem 4.1 and Lemma 4.2.
Thus, let $F\in\mathcal{D}(G_{\gamma})$ and assume in this section that $\tau_{F}>0.$ By the remark below Lemma 4.2, $g(x)=x^{\gamma_{+}},$ where $\gamma_{+}=\max(\gamma,0)$. Moreover, by Potter’s bounds, for arbitrary $\delta_{1},\delta_{2}>0$, there exists $T=T(\delta_{1},\delta_{2})$ such that for $t\geq T$
$g_{1}(x):=\max\big{(}(1-\delta_{1})x^{\gamma_{+}-\delta_{2}},1\big{)}\leq f(x,t)\leq(1+\delta_{1})x^{\gamma_{+}+\delta_{2}}=:g_{2}(x),$
@@
Next, if $\gamma>0$ then we can use the second-order Condition 2. To derive an analogous condition in case $\gamma\leq 0,$ recall that $F\in\mathcal{D}(G_{\gamma})$ is equivalent to
$\lim_{t\to\infty}\frac{U_{F}(tx)-U_{F}(t)}{a(t)}=h_{\gamma}(x),\quad x>0,$ (A.1)
@@
$\lim_{t\to\infty}\frac{U_{F}(tx)/U_{F}(t)-x^{\gamma_{+}}}{a(t)/U_{F}(t)}=h_{\gamma}(x),$ (A.2)
@@
$\tilde{g}(x)=x^{\gamma_{+}}h_{\hat{\rho}}(x)$ (A.3)
with $\hat{\rho}=\rho I(\gamma>0)+\gamma I(\gamma\leq 0)$ and $\hat{a}$ equal to $a$ in Condition 2 for $\gamma>0$ and $a/U$ in (A.2) for $\gamma\leq 0,$ respectively. Next, the conditions of Lemma 4.6, [28], are satisfied under Condition 2 for $\gamma>0$ and (A.2) for $\gamma\leq 0,$ respectively, and thus in (4.12) we can set $\hat{g}(x)=Kx^{\gamma_{+}+\delta}$ for some $K>0,$ $\delta>0$ and $t\geq T.$ Finally, since $\hat{a}$ is regularly varying in our case, it suffices to require $\sqrt{k}\hat{a}(n/k)\to\lambda$ instead of $\sqrt{k}\hat{a}(\xi_{n-k,n})\stackrel{{\scriptstyle P}}{{\to}}\lambda$ by Smirnov’s lemma (see, e.g., Lemma 2.2.3 in [11] or (B.53)).
@@
    Let $F\in\mathcal{D}(G_{\gamma})$ with $\gamma\in\amsmathbb{R}$ and $\tau_{F}>0$, and continuous and eventually monotone $\vartheta:[1,\infty)\to\amsmathbb{R}$ satisfy $\amsmathbb{E}\big{[}|\vartheta(\xi^{\gamma_{+}+\delta})|^{1+\delta}\big{]}<\infty$ for some $\delta>0,$ and $\xi\sim\text{Pareto}(1).$ Assume $\mu(t),$ $k$ and $\xi_{1},\xi_{2},\ldots$ are as in Theorem 4.1. Then $R_{k,n}$ is consistent for $\amsmathbb{E}\big{[}\vartheta(\xi^{\gamma_{+}})\big{]}.$ If, moreover, $\amsmathbb{E}\big{[}|\vartheta(\xi^{\gamma_{+}+\delta})|^{2+\delta}\big{]}<\infty$ for some $\delta>0,$ then
@@
    Assume additionally Condition 2 for $\gamma>0,$ $\sqrt{k}\hat{a}(n/k)\to\lambda$ with finite $\lambda$ for $\hat{a}$ defined above, and $\vartheta^{\prime}$ is continuous and eventually monotone. Finally, assume $\amsmathbb{E}\big{[}\xi^{\gamma_{+}+\delta}|\vartheta^{\prime}(\xi^{\gamma_{+}+\delta})|\big{]}<\infty$ for some $\delta>0.$ Then
@@
    where $C=\amsmathbb{E}\big{[}\tilde{g}(\xi)\vartheta^{\prime}(\xi^{\gamma_{+}})\big{]}$ and $\tilde{g}$ is defined by (A.3).
Remark A.2.
In the above, it is always possible to find an intermediate sequence $k=k_{n}$ such that $\sqrt{k}\hat{a}(n/k)\to\lambda$ is satisfied. Indeed, if $\gamma>0$ then $\hat{a}(t)\to 0$ as $t\to\infty$ by definition (see Condition 2), and if $\gamma\leq 0$ then $\hat{a}(t)\to 0$ as $t\to\infty$ by Lemma 1.2.9, [11].
Note again that to estimate the extreme value index, Corollary A.1 is not enough since for negative $\gamma$, $R_{k,n}$ is a consistent estimator of $\vartheta(1)$, which itself does not depend on $\gamma.$
Consequently, we now consider another candidate function $f(x,t).$ As a starting point, note that it follows from (A.1) that
$\lim_{t\to\infty}\frac{\log U_{F}(tx)-\log U_{F}(t)}{a(t)/U_{F}(t)}=h_{\gamma_{-}}(x)$ (A.4)
with $\gamma_{-}=\min(\gamma,0)$ and the same $a(t)$ as in (A.1); see, e.g., formula (3.5.4) in [11]. The limit (A.4) implies that $\log U_{F}(t)$ is extended regularly varying with index $\gamma_{-},$ see the properties of such functions in section B.2, [11]. The relation (A.4) has previously been used to establish the asymptotic behavior of the moment estimator of EVI; see [12]. Thus, let us consider the candidate function
@@
$x\mapsto\gamma^{-1}\big{(}\max[(1\mp\delta)x^{\gamma\mp\delta}-(1\pm\delta),0]\big{)}$ (A.5)
@@
Recall that the second-order condition for $\log U_{F}(t)$ is given by (4.13). Note that (4.13) follows from the second-order condition, which is standard for extreme value theory,
$\lim_{t\to\infty}\frac{\frac{U_{F}(tx)-U_{F}(t)}{a(t)}-h_{\gamma}(x)}{A(t)}=H_{\gamma,\rho}(x),$ (A.6)
@@
$\left|\frac{\log U_{F}(tx)-\log U_{F}(t)}{a(t)/U_{F}(t)}-h_{\gamma_{-}}(x)\right|\leq K|Q(t)|\,x^{\gamma_{-}+\delta},\quad x\geq 1,t\geq T.$ (A.7)
Clearly, $\delta$ in (A.7) can be chosen such that $\gamma+\delta<0$ for negative $\gamma.$ Finally, $Q(t)$ appearing in (4.13) is regularly varying with index $\rho^{\prime}$ by Theorem B.3.1, [11], therefore as above we can require $\sqrt{k}Q(n/k)\to\lambda$ instead of $\sqrt{k}Q(\xi_{n-k,n})\to\lambda.$
@@
for some continuous function $\vartheta$ and denote
@@
    Let $F\in\mathcal{D}(G_{\gamma})$ with $\gamma\in\amsmathbb{R}$ and $\tau_{F}>0.$ Let $\vartheta:[0,\infty)\to\amsmathbb{R}$ be continuous and eventually monotone, and if $\gamma\geq 0,$ satisfy $|\vartheta(x)|\leq A\max(x^{c},1),$ $x\geq 0,$ for some $A>0$ and $c>0.$ Assume $k$ and $\xi_{1},\xi_{2},\ldots$ are as in Theorem 4.1. Then $T_{k,n}$ is consistent for $\amsmathbb{E}\big{[}\vartheta(h_{\gamma_{-}}(\xi))\big{]}$ and
@@
    Assume additionally (4.13), $\sqrt{k}Q(n/k)\to\lambda$ with finite $\lambda$ for $Q$ defined in (4.13), and continuous and eventually monotone $\vartheta^{\prime}$ is such that $|\vartheta^{\prime}(0)|<\infty.$ Finally, if $\gamma\geq 0,$ assume $|\vartheta^{\prime}(x)|\leq A\max(x^{c},1),$ $x\geq 0,$ for some $A>0$ and $c>0.$ Then
@@
In the same way that Corollary A.1 did not yield a suitable estimators when operating in the $\gamma<0$ domain, Corollary A.3 is again not entirely suitable for all max-domains. Although it succeeds for when $\gamma\leq 0$, the $\gamma>0$ domain fails, since then $T_{k,n}$ converges in probability to $\amsmathbb{E}[\vartheta(\eta)],$ where $\eta\sim\mbox{Exp}(1).$ However, the tools developed in this section are sufficient – upon suitable combination – for the $\gamma\in\amsmathbb{R}$ case.
Appendix B Proofs of Section 3
Notation
First, define the random vectors $\{(V^{t}_{i},\delta^{t}_{i})\}_{i=1}^{k}$ such that
$\bar{D}^{t}_{k}:=\left(\binom{V^{t}_{1,k}}{\delta^{t}_{[1:k]}},\dots,\binom{V^{t}_{k,k}}{\delta^{t}_{[k:k]}}\right)\stackrel{{\scriptstyle d}}{{=}}\left(\binom{Z_{n-k+1,n}/Z_{n-k,n}}{\delta_{[n-k+1:n]}},\ldots,\binom{Z_{n,n}/Z_{n-k,n}}{\delta_{[n:n]}}\right)$ (B.1)
given $Z_{n-k,n}=t,$ $k<n,$ where $\{\delta^{t}_{[i:k]}\}_{i=1}^{k}$ denote concomitants of $\{\delta^{t}_{i}\}_{i=1}^{k}.$
The following notation provides the classical big-o and small-o notation adapted to uniformly control the tail of extreme value estimators. It is used throughout the proofs.
Definition 3.
For the random vectors $\bar{D}^{t}_{k}$ introduced above, an array of functions $(f_{k,n})_{k,n\in\amsmathbb{N}}$, and a positive deterministic sequence $a_{1},a_{2},\ldots$ we say that $f_{k,n}(\bar{D}^{t}_{k})$ is $o_{{P}}(a_{k})$ uniformly for large $t$ if for every $\epsilon>0$,
$\lim_{n\to\infty}\limsup\limits_{t\rightarrow\infty}{\amsmathbb{P}}(|f_{k,n}(\bar{D}^{t}_{k})/a_{k}|>\epsilon)=0.$ (B.2)
We write for short (when no confusion arises) that $f_{k,n}(\bar{D}^{t}_{k})=\overline{o}_{P}(a_{k})$. Analogously $f_{k,n}(\bar{D}^{t}_{k})=\overline{O}_{P}(a_{k})$ using the notions of bounded in probability in place of convergence to zero in probability.
Notice that according to (B.1), $r_{k,n}=r_{k,n}\big{(}(Z^{\ast}_{i},\delta_{i})_{i=1}^{k}\big{)}$ defined in Theorem 3.1 is equal in distribution to $r^{t}_{k,n}:=r_{k,n}(\bar{D}^{t}_{k})$ given $Z_{n-k,n}=t.$ Then, since the latter relation is a distributional property, Theorem 3.1 states that $r^{t}_{k,n}=\overline{o}_{P}(k^{-1/2}).$
Proof of Theorem 3.1.
The following condition serves as a building block towards the general case.
Condition 7.
The function $\varphi$ satisfies
$\varphi(x)=0\>\>\text{ for all }\>\>x>T\>\>\text{ and some }\>\>T<\infty.$ (B.3)
Theorem B.1 (Decomposition, truncated case).
Let $\varphi$ be continuous and satisfy Condition 7. Assume the conditions imposed on $F$ and $G$ in Theorem 3.1. Let $k=k_{n}$ be an intermediate sequence. Then the random sequence $r^{t}_{k,n}$ is $\overline{o}_{P}(k^{-1/2})$.
Proof of Theorem B.1.
We can find i.i.d. random variables $\{V^{t}_{i}\}_{i=1}^{k}$ such that the joint distribution of the set of their order statistics $(V^{t}_{1,k},\ldots,V^{t}_{k,k})$ equals the joint distribution of the set $(Z_{n-k+1,n}/Z_{n-k,n},\ldots,Z_{n,n}/Z_{n-k,n})$ given $Z_{n-k,n}=t,$ which follows by slight modification of lemma 3.4.1 in [11]. But it is not evident whether the random variables $\{\delta^{t}_{i}\}_{i=1}^{k}$ introduced at the beginning of this section are i.i.d. as well. The following lemma answers this question.
Lemma B.2.
The sequence of vectors $\{(V_{i}^{t},\delta^{t}_{i})\}_{i=1}^{k}$ is i.i.d.
Proof.
Since the distribution of $\{(Z_{i},\delta_{i})\}_{i=1}^{n}$ is completely defined by the distribution of $\{(X_{i},Y_{i})\}_{i=1}^{n},$ let us consider the latter. Let $\{T_{i}\}_{i=1}^{k}$ be the times of exceedances over $t$ by the sequence $\{Z_{i}\}_{i=1}^{n},$ that is
$T_{i}=\min\{j>T_{i-1}:\;Z_{j}>t\},\quad T_{0}=0.$
Clearly, the total number of $T_{i}$ is equal to $k$ given $Z_{n-k,n}=t.$ Consider the joint distribution of the pairs $\{(X_{T_{i}},Y_{T_{i}})\}_{i=1}^{k}$ given $Z_{n-k,n}=t.$ First, the condition $Z_{n-k,n}=t$ means that $X_{T_{i}}\geq t,$ $Y_{T_{i}}\geq t,$ $i=1,\ldots,k,$ so it is easy to show that
Clearly, given that the number of exceedances is exactly $k,$ the vector $(T_{1},\ldots,T_{k})$ is uniformly distributed on the set of all k-tuples of $\{1,\ldots,n\},$ thus the latter distribution is equal to
$\mathcal{L}\left(\binom{X^{t}_{1}}{Y^{t}_{1}},\ldots,\binom{X^{t}_{k}}{Y^{t}_{k}}\right),$
where $\{X_{i}^{t}\}_{i=1}^{k}$ and $\{Y_{i}^{t}\}_{i=1}^{k}$ are two independent i.i.d. samples with cdfs $F^{t}(x/t)$ and $G^{t}(x/t),$ respectively. Since the distribution of $(Z_{i},\delta_{i})$ is completely determined by the distribution of $(X_{i},Y_{i}),$ we can select $V_{i}^{t}=\min(X_{i}^{t},Y_{i}^{t})/t$ and $\delta_{i}^{t}=I(X_{i}^{t}\leq Y_{i}^{t}).$
Consider now the extreme Kaplan–Meier integral $S_{k,n}(\varphi),$ defined by (3.3), given $Z_{n-k,n}=t.$ By (B.1) and Lemma B.2, under this condition the extreme Kaplan–Meier estimator (3.2) equals in distribution to
$\amsmathbb{F}_{k,n}^{t}(x)=1-\prod_{i=1}^{k}\left[1-\frac{\delta^{t}_{[i:k]}}{k-i+1}\right]^{I(V^{t}_{i,k}\leq x)},$
and replacing in the latter $k$ by $n,$ $\delta^{t}_{[i:k]}$ by $\delta_{[i:n]},$ and $V_{i,k}^{t}$ by $Z_{i,n}$ gives us exactly the classical Kaplan–Meier estimator (2.2). In particular, this implies that, for fixed $k$ and $n$, we may follow along the lines of the initial steps of the proof of Theorem 1.1 in [32] (specifically, compare the below formula with (2.1) therein) to obtain
$S_{k,n}(\varphi)\,{\stackrel{{\scriptstyle d}}{{=}}}\,\frac{1}{k}\sum_{i=1}^{k}\varphi(V_{i}^{t})\gamma^{t}_{0}(V_{i}^{t})\delta^{t}_{i}[1+B_{ik}^{t}+C_{ik}^{t}]+\frac{1}{k}\sum_{i=1}^{k}\frac{1}{2}\varphi(V_{i}^{t})\delta^{t}_{i}e^{\Delta_{i}^{t}}[B_{ik}^{t}+C_{ik}^{t}]^{2},$ (B.4)
where
$B_{ik}^{t}=k\int_{1}^{V_{i}^{t}-}\log\left[1+\frac{1}{k(1-H_{k}^{t}(z))}\right]H^{0,t}_{k}(\mathop{}\!\mathrm{d}z)-\int_{1}^{V_{i}^{t}-}\frac{1}{1-H_{k}^{t}(z)}H^{0,t}_{k}(\mathop{}\!\mathrm{d}z),$ (B.5)
$C_{ik}^{t}=\int_{1}^{V_{i}^{t}-}\frac{1}{1-H_{k}^{t}(z)}H^{0,t}_{k}(\mathop{}\!\mathrm{d}z)-\int_{1}^{V_{i}^{t}-}\frac{1}{1-H^{t}(z)}H^{0,t}(\mathop{}\!\mathrm{d}z),$ (B.6)
$\Delta_{i}^{t}$ is between the first term in the rhs of (B.5) and second term in the rhs of (B.6), and
$H^{t}_{k}(z)=\frac{1}{k}\sum_{i=1}^{k}I(V_{i}^{t}\leq z),\quad H^{j,t}_{k}(z)=\frac{1}{k}\sum_{i=1}^{k}I(V_{i}^{t}\leq z,\delta_{i}^{t}=j),\quad j=0,1.$
Moreover,
$\frac{1}{k}\sum_{i=1}^{k}\varphi(V_{i}^{t})\gamma^{t}_{0}(V_{i}^{t})\delta^{t}_{i}C_{ik}^{t}=$ $-\int\int\int\frac{I(v<u,v<w)\varphi(w)\gamma_{0}^{t}(w)}{(1-H^{t}(v))^{2}}H_{k}^{t}(\mathop{}\!\mathrm{d}u)H_{k}^{0,t}(\mathop{}\!\mathrm{d}v)H_{k}^{1,t}(\mathop{}\!\mathrm{d}w)$
$+2\int\int I(v<w)\frac{\varphi(w)\gamma_{0}^{t}(w)}{1-H^{t}(v)}H^{0,t}_{k}(\mathop{}\!\mathrm{d}v)H^{1,t}_{k}(\mathop{}\!\mathrm{d}w)$ (B.7)
$-\int\int I(v<w)\frac{\varphi(w)\gamma_{0}^{t}(w)}{1-H^{t}(v)}H^{0,t}(\mathop{}\!\mathrm{d}v)H^{1,t}_{k}(\mathop{}\!\mathrm{d}w)+r_{k,n}^{1,t},$
where
$r_{k,n}^{1,t}=\int\int\varphi(w)\gamma_{0}^{t}(w)I(z<w)\frac{(H_{k}^{t}(z)-H^{t}(z))^{2}}{(1-H^{t}(z))^{2}(1-H^{t}_{k}(z))}H^{0,t}_{k}(\mathop{}\!\mathrm{d}z)H^{1,t}_{k}(\mathop{}\!\mathrm{d}w).$ (B.8)
As a consequence of Condition 7, all denominators appearing in the proof are bounded away from zero. Hereinafter, integrals without integration limits are understood to be from 1 to $\infty$.∞ . The following result is the analogue of Lemma 2.3, [32], in our setting.
Lemma B.3.
Proof.
Consider the second term in the rhs of (B.7)
$V^{t}_{k,n}$ $:=\int\int I(v<w)\frac{\varphi(w)\gamma_{0}^{t}(w)}{1-H^{t}(v)}H^{0,t}_{k}(\mathop{}\!\mathrm{d}v)H^{1,t}_{k}(\mathop{}\!\mathrm{d}w)$
$=\frac{1}{k^{2}}\sum_{i=1}^{k}\sum_{j=1}^{k}\frac{I(V_{i}^{t}<V_{j}^{t})(1-\delta_{i}^{t})\delta_{j}^{t}\varphi(V_{j}^{t})\gamma_{0}^{t}(V_{j}^{t})}{1-H^{t}(V_{i}^{t})}.$
We now rely on several results from classical U-statistics, specifically, Lemma 5.2.1A and Theorem 5.3.2 in [29]. First, note that $V^{t}_{k,n}$ is in fact a V-statistic (see for definition, e.g., Section 5.1.2, [29]). Next, define
$U^{t}_{k,n}:$ $=\frac{k}{k-1}V^{t}_{k,n}$
$=\frac{1}{k(k-1)}\sum_{i=1}^{k}\sum_{j\neq i}\frac{I(V_{i}^{t}<V_{j}^{t})(1-\delta_{i}^{t})\delta_{j}^{t}\varphi(V_{j}^{t})\gamma_{0}^{t}(V_{j}^{t})}{1-H^{t}(V_{i}^{t})}$
$=\frac{1}{k(k-1)}\sum_{i=2}^{k}\sum_{j=1}^{i-1}\left(\frac{I(V_{i}^{t}<V_{j}^{t})(1-\delta_{i}^{t})\delta_{j}^{t}\varphi(V_{j}^{t})\gamma_{0}^{t}(V_{j}^{t})}{1-H^{t}(V_{i}^{t})}\right.$
$\qquad\qquad\qquad\qquad+\left.\frac{I(V_{j}^{t}<V_{i}^{t})(1-\delta_{j}^{t})\delta_{i}^{t}\varphi(V_{i}^{t})\gamma_{0}^{t}(V_{i}^{t})}{1-H^{t}(V_{j}^{t})}\right)$
$=\frac{1}{k(k-1)}\sum_{c}\left(h_{0}\Big{(}\binom{V_{i}^{t}}{\delta_{i}^{t}},\binom{V_{j}^{t}}{\delta^{t}_{j}}\Big{)}+h_{0}\Big{(}\binom{V_{j}^{t}}{\delta^{t}_{j}},\binom{V_{i}^{t}}{\delta_{i}^{t}}\Big{)}\right)$
$=\frac{2}{k(k-1)}\sum_{c}h\Big{(}\binom{V_{i}^{t}}{\delta_{i}^{t}},\binom{V_{j}^{t}}{\delta^{t}_{j}}\Big{)},$
with $\textbf{x},\textbf{y}\in\amsmathbb{R}^{2},$ and $h(\textbf{x},\textbf{y})=\big{[}h_{0}(\textbf{x},\textbf{y})+h_{0}(\textbf{y},\textbf{x})\big{]}/2.$ Clearly, the above expression casts $U^{t}_{k,n}$ as a U-statistic, since the kernel $h$ is symmetric. Denote
$\theta^{t}=\amsmathbb{E}h\Big{(}\binom{V_{1}^{t}}{\delta_{1}^{t}},\binom{V_{2}^{t}}{\delta^{t}_{2}}\Big{)}.$
Then, according to the definition from Section 5.3.1, [29], the Hajek projection of the U-statistic $U^{t}_{k,n}$ is